1.      The demand function for plantains is Qd =100 – 6P while the supply for plantains is

Qs = 28 + 3P where Qd is the quantity demanded, Qs is the quantity supplied and P is the price of plantains.

(a)   Calculate the equilibrium price and quantity for plantains.

(b)   Plot the data on a graph

(c)    On the same graph increase the supply of plantains,.

(d)   State what is the impact on equilibrium price and quantity?


1
Expert's answer
2022-01-20T09:54:07-0500

(a) Qd=Qs

100 – 6P=28 + 3P

72=9P

P=8

"Q=28+3\times8=52"


(b)




(c)



(d)The price will drop. the number will increase